# DriveGPT: Scaling Autoregressive Behavior Models for Driving

### v2

@@ 1 Introduction
Transformer-based foundation models have become increasingly prevalent in sequential modeling tasks across various machine learning domains. These models are highly effective in handling sequential data by capturing long-range dependencies and temporal relationships. Their success has been evident in natural language processing [15], time-series forecasting [30], and speech recognition [13], where sequential patterns play a crucial role. One of the key strengths of transformer-based models is their capacity to learn from large datasets including millions of training examples, enabling them to address complex tasks with increased model sizes, up to billions of model parameters.
@@
VectorNet [4] 72K 211K
PlanTF [1] 2.1M 1M
QCNet-Argo2 [31] 7.3M 200K
QCNet-WOMD [31] 7.5M 2.2M
Wayformer [17] 20M 2.2M
MotionLM [21] 27M 1.1M
MTR [22] 66M 2.2M
GUMP [7] 523M 2.6M
@@
While scaling up model and dataset sizes has been critical for recent advances in sequential modeling for text prediction [11], it remains unclear whether these scaling trends can be directly extended to behavior modeling, particularly in driving tasks, due to several unique challenges. First, driving tasks involve a wider range of input modalities, including agent trajectories and map information, unlike language tasks that rely solely on textual inputs. Second, behavior modeling demands spatial reasoning and an understanding of physical kinematics. Such capabilities are typically beyond the scope of language models. Finally, the collection of large-scale driving datasets requires substantial effort and resources, making it far more challenging than gathering textual data. As a result, existing work is often constrained by the availability of training data or the scalability of the models, as summarized in Table 1.
@@
As we scale up the volume of training data and the number of model parameters, we observe improvements in both quantitative metrics and qualitative behaviors. More importantly, large models trained on extensive, diverse datasets can better handle rare or edge-case scenarios, which often pose significant challenges for autonomous vehicles, as shown in Fig. 1. As a result, we see great potential in scaling up behavior models through data and model parameters to improve the safety and robustness of autonomous driving systems.
@@
    We quantitatively and qualitatively compare models from our scaling experiments to validate their effectiveness in real-world driving scenarios. We present real deployment of our model through closed-loop driving in challenging conditions.
@@ 2.1 Behavior Modeling
Behavior modeling is a critical task in autonomous driving, which covers a broad spectrum of tasks including planning, prediction, and simulation. Taking multimodal inputs including agent history states and map information, behavior models predict the future states of these traffic agents by reasoning about agent dynamics [2, 23], interactions [24, 9], and driving environments [12, 14].
Among learning-based models, transformers have gained popularity due to their ability to fuse multimodal inputs as encoders [17, 29, 31, 8] and model long-range temporal relationships as decoders [21, 22]. Despite the success of transformers in behavior modeling, existing literature is often restricted by the size of model parameters due to limited training data, which fails to capture the full scaling potential of transformer-based models. In our work, we scale up our transformer models to include billions of parameters, by training on a large-scale dataset including more than 100M driving demonstrations, and validate the scalability of transformer-based models in the context of autonomous driving.
@@ 2.2 Large Transformer Models
Large transformers have demonstrated great success in sequential modeling tasks, by scaling up model parameters and data sizes [11, 6, 16, 28]. These scaling laws have pushed the boundary of many sequential modeling tasks including natural language processing [15], time-series forecasting [30], and speech recognition [13].
Recent work has studied the scalability of behavior models in the context of motion prediction [21], planning [25], and simulation [7], yet these studies are either constrained by limited data size (up to a couple of million training examples), or focused on a few orders of magnitude in terms of data and model scaling, limiting the potential to draw statistically significant conclusions over a large scaling range.
In this paper, we study the scaling properties (in terms of data samples, model parameters, and compute) across a much larger range compared to prior work. More specifically, we target an autoregressive decoder architecture that has been proven to be both scalable (as in the LLM literature) and effective in generating accurate trajectories for traffic agents [21].
Beyond autonomous driving, there is limited relevant literature on building large transformer models for robotic tasks [18, 26], which share a transformer architecture similar to our work. Robotics models often share different input features and dynamic models, and operate in different environments, making them difficult to apply directly to driving tasks.
@@ 3.1 Problem Formulation
We define “state” as a full kinematic state including position, orientation, velocity, and acceleration, which is commonly available in agent history observations, and “position” as 2-D (x, y) coordinates to simplify the output space.
@@ 3.2 Scene Encoder
The encoder follows a standard transformer encoder architecture [22] that fuses all input modalities into a set of scene embedding tokens. It consumes raw input features, including target agent history states, nearby agent history states, and map states as a set of vectors, and normalizes all the inputs to an agent-centric view. Each vector is mapped to a token embedding through a PointNet-like encoder as in [4]. At the end of the encoder, we apply a self-attention transformer [27] to fuse all input context into a set of encoder embeddings, $\mathbf{c}\in\mathbb{R}^{n\times d}$, where $n$ is the number of vectors and $d$ is the token dimensions, that summarize the driving scene.
@@ 3.3 LLM-Style Trajectory Decoder
Inspired by the LLM literature [19], we follow [21] to use a transformer decoder architecture to predict the distribution of agent positions at each step in the future.
@@
The output is a set of discrete actions $a$ represented as the Verlet action [20, 21], as the second derivative of positions. We can apply the following equation to map Verlet actions to positions:
@@ 3.4 Training
To train a DriveGPT model, we follow teacher forcing by applying ground truth future positions as inputs to the trajectory decoder. This allows us to predict all future steps in parallel.
@@ 3.5 Inference
We sample multiple trajectories in batch to approximate the distribution and then subsample to the desired number of modes using $K$-Means, as in [21].
@@ 4 Scaling Experiments
The goal of our scaling experiments is to determine the effect of data and model size on behavior prediction performance. Quantifying scaling laws similar to those seen in LLMs [11] can help prioritize the value of data and compute for future research directions in behavior modeling. We focus our effort on exploring the next frontier of data and model size — over an order of magnitude beyond previously published work.
4.1 Large-Scale Driving Dataset
From millions of miles of high-quality real-world human driving demonstrations, we curated a small subset of 120M segments for an internal research dataset. The dataset is carefully curated and balanced to cover diverse geographic regions from multiple cities and countries, and a variety of challenging driving scenarios, such as passing a double-parked vehicle, unprotected left turns, construction zones, etc. We extracted map information, target agent states, and nearby agent states into vectorized representation, as customary in behavior modeling literature [4].
We scale the model size across three orders of magnitude, from 1.5 million to 1.4 billion parameters, by increasing the embedding dimension. For each model size, we explore multiple learning rate schedules and select the one that yields the best performance. Consistent with practices in large language model scaling [11], each model is trained for a single epoch.
We measure model performance through validation loss computed over a comprehensive validation set. We use validation loss as a proxy to measure model performance, following standard practices in scaling studies [11, 6, 16]. This loss, calculated as cross-entropy on next-action prediction, serves as our primary performance metric. Additional driving-specific metrics are reported in Sec. 5.1.1 and Sec. 5.1.2.
@@ 4.2 Data Scaling
Data scaling results are summarized in Fig. 3. The smallest dataset of 2.2M samples mimics the size of Waymo Open Motion Dataset (WOMD) [3], a large open-source dataset for behavior modeling ($\sim$44k scenarios with multiple target agents per scenario). We select a few subsets of our internal research dataset to study data scaling across different orders of dataset sizes. Our experiments use $\sim$50x more data than WOMD, exploring a new region of the design space.
@@ 4.3 Model Scaling
We now study model sizes across three orders of magnitude (1.5M to 1.4B parameters). We increase model size by increasing the hidden dimensions of transformers for simplicity. We notice that modifying other parameters such as number of attention heads and hidden dimensions per head does not lead to noticeable changes in the results, as studied in Appendix B.
Training larger models is sensitive to learning rates, as observed in other scaling studies [11, 6]. For each model size, we run multiple experiments at different learning rates to select the one with the optimal performance, as summarized in Appendix B.
@@
Results in Fig. 4 demonstrate that increasing the amount of training data enhances the effectiveness of model scaling. Specifically, when the dataset size is up to 21M, the impact on the validation loss is barely noticeable across a large range of model sizes. Beyond 21M samples, validation losses improve with larger models – up to 12M parameters when trained on the 42M dataset and up to 94M parameters when trained on the 120M dataset – before reaching a plateau and eventually overfitting. These findings further reinforce that data is the primary bottleneck for scaling models, aligning with observations in the LLM scaling literature [11].
4.4 Compute Scaling
In Fig. 5, we examine how compute affects training loss, where compute is measured by Floating Point Operations (FLOPs). We identify a monotonically decreasing “min-bound” boundary, which shows the lowest training loss observed up to the current compute value. As we increase compute, training loss generally decreases. Initially, this decrease is quite steep, but it gradually slows down at higher FLOPs values. This trend is consistent with observations in the LLM scaling literature, such as those reported by [6], covering a subset of the full FLOP range explored in these studies.
Next, we investigate if there is an optimal combination of data size and model parameters for a fixed compute budget. Given the large computational expense for training models at the scales we are exploring, it is important to make the best use of our data. In this study, we fixed the compute budget at different FLOP groups. For a fixed compute budget, we can allocate resources either to model parameters or data samples, keeping their product constant.
Fig. 6 plots performance with different compute budgets. The trend clearly shows that a larger compute budget leads to better performance, with optimal model size increasing accordingly, as indicated by the “best” gray line. The results further reveal that data is the main bottleneck, as the smallest model outperforms others in the three largest FLOP groups.
@@ 4.5 Ablation Study on Decoder Architecture
We scale up two different model architectures by two orders of magnitude: our autoregressive decoder and a one-shot decoder. For the one-shot decoder, we follow [17] to use a transformer decoder that takes a set of learned queries and cross attends them with scene embeddings to produce trajectory samples. This decoder is referred to as “one-shot” because it generates the full trajectory rollout at once, where an autoregressive decoder follows an LLM-style to produce trajectories one step at a time.
@@ 5.1 Internal Evaluation: AV Planning
We approximate the distribution by oversampling trajectories in batch and subsampling to 6 trajectories as in [21]. While the AV must ultimately select a single trajectory for planning, returning multiple samples helps better understand multimodal behavior and aligns with motion prediction metrics.
We measure the planning performance on a comprehensive test set through a set of standard geometric metrics including minADE (mADE), minFDE (mFDE), and miss rate (MR). Additionally, we use semantic-based metrics including offroad rate (Offroad) that measures the ratio of trajectories that leave the road and collision rate (Collision) that measures the ratio of trajectories overlapping with static agents. We normalize these metrics across experiments to highlight relative performance changes.
@@
Figure 8: By training on 50x more data, DriveGPT is able to produce high-quality trajectories in red that keep a safe lateral distance from a jaywalker (top) and go around two double-parked vehicles (bottom).
@@ 5.1.1 Data Scaling Results
The results are presented in Table 2, where we see that training on more data samples significantly improves the quality of predicted AV trajectories, in terms of critical semantics metrics in driving including offroad rate and collision rate, as well as geometric metrics. These improvements are consistent with Sec. 4.2.
@@ 5.1.2 Model Scaling Results
The results are presented in Table 6, where all metrics improve as the model size increases up to 94M parameters, with further gains in the collision metric at 163M parameters. Although validation loss plateaus at 94M in Fig. 4, driving metrics continue to improve with increased model capacities, highlighting the potential benefits of introducing more parameters for enhanced driving performance.

### v3

@@ 1 Introduction
Transformer-based foundation models have become increasingly prevalent in sequential modeling tasks across various machine learning domains. These models are highly effective in handling sequential data by capturing long-range dependencies and temporal relationships. Their success has been evident in natural language processing (Mann et al., 2020; Kaplan et al., 2020; Hoffmann et al., 2022), time-series forecasting (Zhou et al., 2021), and speech recognition (Kim et al., 2022), where sequential patterns play a crucial role. One of the key strengths of transformer-based models is their capacity to learn from large datasets including millions of training examples, enabling them to address complex tasks with increased model sizes, up to billions of model parameters.
@@
VectorNet 72K 211K
PlanTF 2.1M 1M
QCNet-Argo2 7.3M 200K
QCNet-WOMD 7.5M 2.2M
Wayformer 20M 2.2M
MotionLM 27M 1.1M
MTR 66M 2.2M
GUMP 523M 2.6M
@@
While scaling up model and dataset sizes has been critical for recent advances in sequential modeling for text prediction (Kaplan et al., 2020; Hoffmann et al., 2022), it remains unclear whether these scaling trends can be directly extended to behavior modeling, particularly in driving tasks, due to several unique challenges. First, driving tasks involve a wider range of input modalities, including agent trajectories and map information, unlike language tasks that rely solely on textual inputs. Second, behavior modeling demands spatial reasoning and an understanding of physical kinematics. Such capabilities are typically beyond the scope of language models. Finally, the collection of large-scale driving datasets requires substantial effort and resources, making it far more challenging than gathering textual data. As a result, existing work is often constrained by the availability of training data or the scalability of the models, as summarized in Table 1.
@@
As we scale up the volume of training data and the number of model parameters, we observe improvements in both quantitative metrics and qualitative behaviors. More importantly, large models trained on extensive, diverse datasets can better handle rare or edge-case scenarios, which often pose significant challenges for autonomous vehicles, as shown in Fig. $1$The baseline model is trained on $\sim$50 times less data and uses $\sim$50 times fewer parameters.. As a result, we see great potential in scaling up behavior models through data and model parameters to improve the safety and robustness of autonomous driving systems.
@@
    We quantitatively and qualitatively compare models from our scaling experiments to validate their effectiveness in real-world driving scenarios. We present real-world deployment of our model through closed-loop driving in challenging conditions.
@@ 2.1 Behavior Modeling
Behavior modeling is a critical task in autonomous driving, which covers a broad spectrum of tasks including planning, prediction, and simulation. Taking multimodal inputs including agent history states and map information, behavior models predict the future states of these traffic agents by reasoning about agent dynamics (Cui et al., 2020; Song et al., 2022), interactions (Sun et al., 2022; Jiang et al., 2023), human intent (Shi et al., 2022; Huang et al., 2022; Sun et al., 2024), and driving environments (Liang et al., 2020; Kim et al., 2021).
Among learning-based models, transformers have gained popularity due to their ability to fuse multimodal inputs as encoders (Nayakanti et al., 2023; Zhang et al., 2023; Zhou et al., 2023; Jia et al., 2023; Gan et al., 2024) and model long-range temporal relationships as decoders (Seff et al., 2023; Shi et al., 2022, 2024). Despite the success of transformers in behavior modeling, existing literature is often restricted by the size of model parameters due to limited training data, which fails to capture the full scaling potential of transformer-based models. In our work, we scale up our transformer models to include billions of parameters, by training on a large-scale dataset including more than 100M driving demonstrations, and validate the scalability of transformer-based models in the context of autonomous driving.
@@ 2.2 Large Transformer Models
Large transformers have demonstrated great success in sequential modeling tasks, by scaling up model parameters and data sizes (Kaplan et al., 2020; Hoffmann et al., 2022; Zhai et al., 2022; Muennighoff et al., 2023). These scaling laws have pushed the boundary of many sequential modeling tasks including natural language processing (Mann et al., 2020), time-series forecasting (Zhou et al., 2021), and speech recognition (Kim et al., 2022).
Recent work has studied the scalability of behavior models in the context of motion prediction (Seff et al., 2023; Ettinger et al., 2024), planning (Sun et al., 2023), and simulation (Hu et al., 2024b), yet these studies are either constrained by limited data size (up to a couple of million training examples), or focused on a few orders of magnitude in terms of data and model scaling, limiting the potential to draw statistically significant conclusions over a large scaling range.
In this paper, we study the scaling properties (in terms of data samples, model parameters, and compute) across a much larger range compared to prior work. More specifically, we target an autoregressive decoder architecture that has been proven to be both scalable (as in the LLM literature) and effective in generating accurate trajectories for traffic agents (Seff et al., 2023; Chen et al., 2024; Hu et al., 2024a).
Beyond autonomous driving, there is limited relevant literature on building large transformer models for robotic tasks (O’Neill et al., 2024; Octo Model Team et al., 2024), which share a transformer architecture similar to our work. Robotics models often share different input features and dynamic models, and operate in different environments, making them difficult to apply directly to driving tasks.
@@ 3.1 Problem Formulation
We define “state” as a complete kinematic state including position, orientation, velocity, and acceleration, which is commonly available in agent history observations, and “position” as 2-D (x, y) coordinates to simplify the output space.
@@ 3.2 Scene Encoder
The encoder follows a standard transformer encoder architecture (Shi et al., 2022) that fuses all input modalities into a set of scene embedding tokens. It consumes raw input features, including target agent history states, nearby agent history states, and map states as a set of vectors, and normalizes all inputs to an agent-centric view. Each vector is mapped to a token embedding through a PointNet-like encoder as in (Gao et al., 2020). At the end of the encoder, we apply a self-attention transformer (Vaswani et al., 2017) to fuse all input context into a set of encoder embeddings, $\mathbf{c}\in\mathbb{R}^{n\times d}$, where $n$ is the number of vectors and $d$ is the token dimensions, that summarize the driving scene.
@@ 3.3 LLM-Style Trajectory Decoder
Inspired by the LLM literature (Radford et al., 2019), we follow (Seff et al., 2023) to use a transformer decoder architecture to predict the distribution of agent positions at each step in the future.
@@
The output is a set of discrete actions $a$ represented as the Verlet action (Rhinehart et al., 2018; Seff et al., 2023), as the second derivative of positions. We can apply the following equation to map Verlet actions to positions:
@@ 3.4 Training
To train a DriveGPT model, we follow teacher forcing by applying ground truth future positions as input to the trajectory decoder. This allows us to predict all future steps in parallel.
@@ 3.5 Inference
We sample multiple trajectories in batch to approximate the distribution and then subsample to the desired number of modes using $K$-Means, as in (Seff et al., 2023).
@@ 4 Scaling Experiments
The goal of our scaling experiments is to determine the effect of data and model size on behavior prediction performance. Quantifying scaling laws similar to those seen in LLMs (Kaplan et al., 2020) can help prioritize the value of data and compute for future research directions in behavior modeling. We focus our effort on exploring the next frontier of data and model size – over an order of magnitude beyond previously published work.
Large-scale driving dataset
From millions of miles of high-quality real-world human driving demonstrations, we curate a small subset of 120M segments for an internal research dataset. The dataset is carefully curated and balanced to represent diverse geographic regions across multiple cities and countries, including the United States, Japan, and the UAE. Data collection is evenly distributed between daytime and nighttime and is conducted primarily in urban environments. The dataset captures a wide range of challenging driving scenarios, such as lane changes, intersections, double-parked vehicles, construction zones, and close interactions with pedestrians and cyclists.
We extracted map information, target agent states, and nearby agent states into vectorized representation, as customary in behavior modeling literature (Gao et al., 2020).
Scaling overview
We scale the model size across three orders of magnitude, from 1.5 million to 1.4 billion parameters, by increasing the embedding dimension in both encoder and decoder transformers. For each model size, we explore multiple learning rate schedules using a cosine decay over the full training steps and select the maximum learning rate that yields the best performance. Table. 2 summarizes the optimal learning rate for each model size. Consistent with practices in large language model scaling (Kaplan et al., 2020), each model is trained for a single epoch.
Model Size Hidden Dimension ($d_{\text{model}}$) Max LR
1.5M 64 0.0050
4M 128 0.0020
8M 192 0.0014
12M 256 0.0010
26M 384 0.0010
94M 768 0.0007
163M 1024 0.0004
358M 1536 0.0002
629M 2048 0.0001
1.4B 3072 0.0001
Table 2: We vary the model size over three orders of magnitude through hidden dimensions. For each model size, we report the optimal learning rate, which decreases as the model size increases, matching the observations in the LLM scaling literature (Kaplan et al., 2020; Hoffmann et al., 2022).
We evaluate model performance using validation loss, computed on a comprehensive validation set of 10 million samples drawn from the same distribution as the training data, with no overlap. This set remains fixed across all scaling experiments to ensure consistency. We use validation loss as a proxy to measure model performance, following standard practices in scaling studies (Kaplan et al., 2020; Hoffmann et al., 2022; Muennighoff et al., 2023). This loss, calculated as cross-entropy on next-action prediction, serves as our primary performance metric. Additional driving-specific metrics are reported in Sec. 5.1.
@@ 4.1 Data Scaling
Data scaling results are summarized in Fig. 3. The smallest dataset of 2.2M samples mimics the size of Waymo Open Motion Dataset (WOMD) (Ettinger et al., 2021), a large open-source dataset for behavior modeling ($\sim$44k scenarios with multiple target agents per scenario). We select a few subsets of our internal research dataset to study data scaling across different orders of dataset sizes. Our experiments use $\sim$50x more data than WOMD, exploring a new region of the design space.
@@ 4.2 Model Scaling
We now study model sizes across three orders of magnitude (1.5M to 1.4B parameters), as listed in Table 2. We increase model size by increasing the hidden dimensions of transformers for simplicity. We notice that modifying other parameters such as number of attention heads and hidden dimensions per head does not lead to noticeable changes in the results, as studied in Appendix B.
Training larger models is sensitive to learning rates, as observed in other scaling studies (Kaplan et al., 2020; Hoffmann et al., 2022). For each model size, we run multiple experiments at different learning rates to select the one with the optimal performance, as summarized in Table. 2.
Results in Fig. 4 demonstrate that increasing the amount of training data enhances the effectiveness of model scaling. Specifically, when the dataset size is up to 21M, the impact on the validation loss is barely noticeable across a large range of model sizes. Beyond 21M samples, validation losses improve with larger models – up to 12M parameters when trained on the 42M dataset and up to 94M parameters when trained on the 120M dataset – before reaching a plateau and eventually overfitting. These findings further reinforce that data is the primary bottleneck for scaling models, aligning with observations in the LLM scaling literature (Kaplan et al., 2020). We leave further exploration of scaling behavior with larger datasets to future work.
@@
4.3 Compute Scaling
In Fig. 5, we examine how compute affects training loss, where compute is measured by Floating Point Operations (FLOPs). We identify a monotonically decreasing “min-bound” boundary, which shows the lowest training loss observed up to the current compute value. As we increase compute, training loss generally decreases. Initially, this decrease is quite steep, but it gradually slows down at higher FLOPs values. This trend is consistent with observations in the LLM scaling literature, such as those reported by (Hoffmann et al., 2022), covering a subset of the full FLOP range explored in these studies.
Next, we investigate whether there is an optimal combination of data size and model parameters for a fixed compute budget. Given the large computational expense for training models at the scales we are exploring, it is important to make the best use of our data. In this study, we fixed the compute budget in different FLOP groups. For a fixed compute budget, we can allocate resources either to model parameters or data samples, keeping their product constant.
Fig. 6 plots the performance with different compute budgets. The trend clearly shows that a larger compute budget leads to better performance, with optimal model size increasing accordingly, as indicated by the “best” gray line. The results further reveal that data is the main bottleneck, as the smallest model outperforms others in the three largest FLOP groups.
@@ 4.4 Ablation Study on Decoder Architecture
We scale up two different model architectures by two orders of magnitude: our autoregressive decoder and a one-shot decoder. For the one-shot decoder, we follow (Nayakanti et al., 2023) to use a transformer decoder that takes a set of learned queries and cross attends them with scene embeddings to produce trajectory samples. This decoder is referred to as “one-shot” because it generates the full trajectory rollout at once, where an autoregressive decoder follows an LLM-style to produce trajectories one step at a time.
@@ 5.1 Internal Evaluation: AV Planning
We approximate the distribution by oversampling trajectories in batch and subsampling to 6 trajectories as in (Seff et al., 2023). While the AV must ultimately select a single trajectory for planning, returning multiple samples helps better understand multimodal behavior and aligns with motion prediction metrics.
We measure the planning performance on a comprehensive test set through a set of standard geometric metrics including minADE (mADE), minFDE (mFDE), and miss rate (MR). Additionally, we use semantic-based metrics including offroad rate (Offroad) that measures the ratio of trajectories that leave the road and collision rate (Collision) that measures the ratio of trajectories overlapping with traffic agents. We normalize these metrics across experiments to highlight relative performance changes.
@@
Figure 8: By training on 50x more data, DriveGPT (red) is able to produce high-quality trajectories in red that keep a safe lateral distance from a jaywalker (top) and go around two double-parked vehicles (bottom).
@@ 5.1.1 Data Scaling Results
The results are presented in Table 3, where we see that training on more data samples significantly improves the quality of the predicted AV trajectories, in terms of critical semantics metrics in driving including offroad rate and collision rate, as well as geometric metrics. These improvements are consistent with Sec. 4.1.
@@ 5.1.2 Model Scaling Results
The results are presented in Table 4, where all metrics improve as the model size increases up to 94M parameters, with further gains in the collision metric at 163M parameters (see Sec. D.1 for qualitative examples). Although validation loss shows diminishing returns beyond 94M in Fig. 4, driving metrics continue to improve with increased model capacities, highlighting the potential benefits of introducing more parameters for enhanced driving performance.